# Quasirandom and quasisimple groups

### v1

@@ 1. Introduction
It is well known that the finite simple groups of Lie type are $\varepsilon$-quasirandom for some $\varepsilon>0$ which depends only on the rank [14], a fact that is exploited in all the papers we mentioned above. A more general discussion on quasirandomness is taken by Babai, Nikolov and Pyber [2], and Nikolov and Pyber [18]. The celebrated Gower’s trick [18] says that, if $G$ is $\varepsilon$-quasirandom and $S\subseteq G$ is a large subset, namely $|S|\geq|G|^{1-\frac{\varepsilon}{3}}$, then $S^{3}=G$. This fact, as well as other natural invariants of $G$, can be used to characterize quasirandomness. Such equivalences look particularly nice when expressed in the exponential regime (see Theorem 2.6 below).
The algebraic structure of quasirandom groups remains unexplored. The present article was born from our need to answer the question: what, concretely, is a quasirandom group? In the first main result, we prove that large $\varepsilon$-quasirandom groups with a trivial solvable radical are direct products of nonabelian simple groups. We say a little more.
@@
Fix $\varepsilon>0$, and let $G$ be an $\varepsilon$-quasirandom group. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/Sol(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
For the second main result, we study the groups that are $\varepsilon$-quasirandom with a relatively large $\varepsilon$. It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the most quasirandom family of groups [24, 29], but not much is known. Not only this is actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
@@ 3. $(c)$
The irreducible representations of the quasisimple groups are known [27], thus it is an easy (although laborious) task to recover the full list of the $\frac{1}{5}$-quasirandom groups from Theorem B. An interesting consequence is the following:
@@
The constant $\frac{3}{14}$ is best possible, precisely because of the family in Theorem Bb. On the other hand, the value $\frac{1}{5}$ is the result of a compromise, as with our machinery and more work one can obtain a deeper classification. There are some reasons to prefer $\frac{1}{5}$: first, it is small enough to provide Corollary C. Second, it makes the theorem sharp in some sense, because there exist infinitely many nonquasisimple groups which are $(\frac{1}{5}-\varepsilon)$-quasirandom, for any $\varepsilon>0$ (see Section 3). It also makes the proof a little cleaner, because every $\frac{1}{5}$-quasirandom group has a unique maximal normal subgroup (Lemma 4.7). We point out that similar methods are useful to study quasirandomnes at other regimes different from the exponential, which could be of interest as well. The shallow version of Theorem B is the following:
@@
We will actually prove Corollary D as a “baby” Theorem B. The philosophy behind the proofs is that quasirandom groups are pretty rigid, and lend themselves very little to nontrivial extensions, especially the noncentral ones. To make this idea really say something, we use various properties of the finite simple groups and their representations, both complex and modular, linear and projective.
The paper is structured as follows. In Section 2 we introduce the (exponential) quasirandom degree $\bm{\varepsilon}(G)$ of a finite group, and give basic results about quasisimple groups, modular and projective representations. In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem Bb. We prove the main theorems in Section 4. Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups: we collect these data in the Appendix.
@@ 2.1. Minimal degree of a nontrivial representation
We start with the following, which appears as [26, Ex. 1.3.2].
@@ 3. $(iii)$
We would like to mention some connections between $\mathbf{D}_{0}(G)$ and other important parameters associated to a finite group: the number of conjugacy classes $\mathbf{k}(G)$, the representation growth $\mathbf{r}_{n}(G)$, and the minimal index of a proper subgroup, namely $\bm{\iota}(G):=\min_{H<G}|G:H|$.
@@
It follows that, if $G$ is an $\varepsilon$-quasirandom group, then $\mathbf{k}(G)\leq|G|^{1-2\varepsilon}$. Moreover, the number $\mathbf{r}_{n}(G)$ of the inequivalent irreducible representations of $G$ of degree $n$ is polynomially bounded, i.e. $\mathbf{r}_{n}(G)\leq n^{\varepsilon^{-1}-2}$ for all $n$ (see [22, Lemma 25]).
@@
for every finite perfect group $G$. A converse inequality, $\mathbf{D}_{0}(G)<\bm{\iota}(G)$, follows from the previous lemmas. Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$). In particular, $\mathbf{D}_{0}(G)<|G|^{0.5}$. As consequences of Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
@@ 2.2. Quasirandom degree
By definition, $\mathbf{D}_{0}(G)=|G|^{\bm{\varepsilon}(G)}$, and $\bm{\varepsilon}(G)$ is the maximum $\varepsilon\in[0,1)$ such that $G$ is an $\varepsilon$-quasirandom group. Observe that $\bm{\varepsilon}(G)>0$ if and only if $G$ is perfect. We now report the theorem of Nikolov and Pyber mentioned in the Introduction:
@@
Let $N\lhd G$ be a nontrivial noncentral abelian normal subgroup. Then
@@ 2.3. Quasisimple groups
It is a key property of perfect groups that, if $G$ is perfect, then there is a unique universal central perfect extension, which we denote by $\widetilde{G}$. This extension is obtained by choosing $N$ isomorphic to the Schur multiplier of $G$, that is, the second homology group $H^{2}(G,\mathbb{Z})$. For more details on this, we refer to [1, Sec. 33].
@@ 2.4. Linear and projective representations
So far, we have only dealt with complex linear representations, but modular and projective representations will be a key tool in the proof of the main theorems. For a prime $r\geq 2$, we use $\mathbf{D}_{r}(G)$ to denote the minimal degree of a nontrivial irreducible modular representation of $G$ over a field of characteristic $r$. We recall that a projective representation of $G$ is a representation of a perfect central extension of $G$. Equivalently, it is a group homomorphism
@@
The degree of the representation is the dimension of $V$. As for linear representations, we denote by $\mathbf{R}_{0}(G)$ the minimal degree of a nontrivial projective complex representation, and by $\mathbf{R}_{r}(G)$ the minimal degree of a nontrivial projective representation over a field of characteristic $r$. We remark that, if $G$ is finite and perfect, and $\widetilde{G}$ is its universal perfect central extension, then
@@
In the remainder of this subsection, we focus on the projective representations of finite simple groups of Lie type. We recall that, if $G$ is a group of Lie type defined over a field of characteristic $p$, then a modular representation of $G$ is in natural characteristic when the characteristic of the underlying vector space is $p$, while it is in cross-characteristic otherwise.
@@
It follows from Propositions 5.3.1, 5.4.6, and Remark 5.4.7 of [13]. ∎
For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom simple groups. The infinite families of $\frac{1}{4}$-quasirandom groups are
@@
The relevant informations can be recovered from [14, 25] for classical groups, and from [15, Sec. 4] for exceptional groups. We collect these data in Table 1.
@@
Moreover, we point out that the comparison between $\mathbf{R}_{0}(L)$ and $\mathbf{R}_{r}(L)$ can be explicitly found, for all simple groups of Lie type, in [27, Sec. 3].
@@ 3. Quasirandom groups of affine type
In this short section, we provide some highly quasirandom nonquasisimple groups. The key tool is Lemma 3.2 below. Before giving the results, we revise some notions from the representation theory of abelian groups. Let $V$ be a finite vector space, and let $V^{*}$ be its dual space. We remark that $V^{*}\cong Hom(V,\mathbb{C}^{*})$. If $\phi:V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}:V^{*}\rightarrow V^{*}$ is defined by
@@
for every $g\in K$.
@@
Let $V=(\mathbb{F}_{q})^{n}$ and $g\in K$. The number of $g$-fixed points in $V$ is $q^{\dim(Ker(g-I))}$, while the number of $g$-fixed points in $V^{*}$ is $q^{\dim(Ker((g^{-1})^{T}-I))}$, where $g^{T}$ is the transpose of $g$. But $\dim(Ker(g-I))$ is the multiplicity of $1$ as an eigenvalue of $g$, so the two quantities above are equal. Since $g$ is arbitrary, the proof follows from Burnside’s lemma. ∎
@@ 4.1. Proof of Theorem A
The main idea behind the proof of Theorem A is the following. Under the hypothesis that $Sol(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$. If this is not the case, then $G/F^{*}(G)$ has a nontrivial simple quotient $L$. We use the quasirandomness of $L$ to obtain an upper bound on $|G|$, while Bender’s theorem and other considerations provide a lower bound on $|F^{*}(G)|$. These inequalities fight each other, giving the desired contradiction. A similar argument will be used in a portion of the proof of Theorem B. Going into details, we start reporting the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
@@
In the following remark, we discuss the asymptotic behaviour of $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$, when $L$ is a finite simple group, and $|L|\rightarrow+\infty$.
@@
Fix $\varepsilon>0$. From Corollary 2.3(i), it is clear that if $|G|$ is large then $|G/Sol(G)|$ is large. Moreover, $G/Sol(G)$ is again $\varepsilon$-quasirandom. Up to replacing $G$ with $G/Sol(G)$, we can assume that $G$ has a trivial solvable radical. Now it is well known that the generalized Fitting subgroup $F=F^{*}(G)$ coincides with the socle of $G$, and it is the direct product of nonabelian simple groups $T_{i}$’s. Let $F=\prod_{i}(T_{i})^{n_{i}}$, where $n_{i}\geq 1$ for each $i$. Moreover, as $\mathbf{C}_{G}(F)\subseteq F$ [1, (31.13)], we have $\mathbf{C}_{G}(F)=\mathbf{Z}(F)=1$. It follows that $G$ can be embedded as a subgroup of
@@
In particular, all the composition factors of $G/F$ lie as sections of one among $\mathrm{Out}(T_{i})$ or $\mathrm{Sym}(n_{i})$, for some $i$. We claim that $G=F$. Otherwise, it is easy to see that $G/F$ has a simple quotient $L$ which is $\varepsilon$-quasirandom. Of course, $|G|\leq\mathbf{D}_{0}(L)^{1/\varepsilon}$. So, if $|L|$ is bounded, then $|G|$ is bounded. It follows that $L=L_{d}(q)$ is a simple group of Lie type, $d$ is $\varepsilon$-bounded, while $q$ is unbounded. By the definition of quasirandom degree, we have
(4.1) $|G|\>\leq\>\mathbf{D}_{0}(L_{d}(q))^{1/\varepsilon}.$
Let $n=\max_{i}n_{i}$, so that $|F|\geq 60^{n}$. From Remark 4.2 and Theorem 4.1, $L_{d}(q)$ has to appear as a section of $\mathrm{Alt}(n)$. From Lemma 4.3, $\mathrm{Sym}(n)$ has a subgroup isomorphic to $L_{d}(q)$. From Lemma 2.2 we obtain $\mathbf{D}_{0}(L_{d}(q))\leq n$, and so
@@
When $\mathbf{D}_{0}(L_{d}(q))$ is sufficiently large, this gives a contradiction to (4.1). But $\mathbf{D}_{0}(L_{d}(q))\geq|G|^{\varepsilon}$, so we obtain a contradiction when $|G|$ is sufficiently large.
We have proven that $G/Sol(G)=F^{*}(G/Sol(G))$, so $G/Sol(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/Sol(G)=\prod_{i}(T_{i})^{n_{i}}$. From Remark 4.4, the $T_{i}$’s are all simple groups of Lie type. Let $N=\sum_{i}n_{i}$. From Corollary 2.9, we have $\varepsilon\leq\mathbf{D}_{0}(G/Sol(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$. Moreover, each block $(T_{i})^{n_{i}}$ has to be $\varepsilon$-quasirandom as a quotient of $G$. From Corollary 2.9 and Remark 4.4, we have
@@
Fix $\varepsilon>0$. After Theorem A, it is natural to ask whether the fact that $G/Sol(G)$ is the direct product of boundedly many simple groups of Lie type of bounded rank, has some effect on the quasirandomness of the full $G$. In general, the answer is no, as it can happen that $G$ is perfect but its quasirandom degree vanishes when $|G|$ grows. It is enough to consider the wreath product $(C_{n})^{5}\rtimes\mathrm{Alt}(5)$ where $n$ is large: here $G/Sol(G)\cong\mathrm{Alt}(5)$ is a perfect group of bounded cardinality, so trivially quasirandom, while the small representations induced from $\mathrm{Alt}(5)$ destroy the quasirandomness of $G$. The following example, where $|G/Sol(G)|$ is unbounded, is due to S. Eberhard.
@@
where $g\in\mathrm{SL}_{2}(p)$, and *** is arbitrary. We have
@@
It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix. In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $\frac{1}{4}$-quasirandom. On the other hand,
$\mathbf{D}_{0}(G_{d})\leq\mathbf{D}_{0}(\mathrm{PSL}_{2}(p))\leq p.$
It follows that $\bm{\varepsilon}(G_{d})\leq d^{-2}$, which vanishes when $d$ grows.
Despite the previous examples, a lower bound for $\bm{\varepsilon}(G)$ as a function of $\bm{\varepsilon}(G/Sol(G))$ exists when $|G/Sol(G)|$ is relatively large.
Lemma 4.6.
Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group. Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $\frac{\log|G/Sol(G)|}{\log|G|}\geq\delta$. If $|G|$ is sufficiently large, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
Proof.
By [8, Lemma 4.10], we have
$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
[…]
and passing to the logarithms,
The proof follows. ∎

### v2

@@ 1. Introduction
It is well known that the finite simple groups of Lie type are $\varepsilon$-quasirandom for some $\varepsilon>0$ which depends only on the rank [15], a fact that is exploited in all the papers we mentioned above. A more general discussion on quasirandomness is taken by Babai, Nikolov and Pyber [2], and Nikolov and Pyber [19]. The celebrated Gowers’ trick [19] says that, if $G$ is $\varepsilon$-quasirandom and $S\subseteq G$ is a large subset, namely $|S|\geq|G|^{1-\frac{\varepsilon}{3}}$, then $S^{3}=G$. This fact, as well as other natural invariants of $G$, can be used to characterize quasirandomness. Such equivalences look particularly nice when expressed in the exponential regime (see Theorem 2.6 below).
The algebraic structure of quasirandom groups remains unexplored. The present article was born from our need to answer the question: what, concretely, is a quasirandom group? In the first main result we describe the large $\varepsilon$-quasirandom groups, for any fixed $\varepsilon>0$. They are precisely the groups whose quotient by the solvable radical is a relatively large direct product of boundedly many finite simple groups of Lie type of bounded rank.
@@
Fix $\varepsilon>0$. If $G$ is an $\varepsilon$-quasirandom group, then $|G/Sol(G)|\geq|G|^{\varepsilon}$. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/Sol(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
Conversely, suppose that $G$ is perfect, $|G/Sol(G)|\geq|G|^{\varepsilon}$, and that $G/Sol(G)$ is $\varepsilon$-quasirandom. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G$ is $\frac{\varepsilon^{2}}{3}$-quasirandom.
For the second main result, we study the groups that are $\varepsilon$-quasirandom with a relatively large $\varepsilon$. It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the “most quasirandom” family of groups [23, 28], but not much is known. Not only this is actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
@@ 3. $(c)$
The low dimensional representations of the quasisimple groups are known [26], thus it is an easy (although laborious) task to recover the full list of the $\frac{1}{5}$-quasirandom groups from Theorem B. An interesting consequence is the following:
@@
The constant $\frac{3}{14}$ is best possible, precisely because of the family in Theorem Bb. On the other hand, the value $\frac{1}{5}$ is the result of a compromise, as with our machinery and more work one can obtain a deeper classification. There are some reasons to prefer $\frac{1}{5}$: first, it is small enough to provide Corollary C. Second, it makes the theorem sharp in some sense, because there exist infinitely many nonquasisimple groups which are $(\frac{1}{5}-\varepsilon)$-quasirandom, for any $\varepsilon>0$ (see Section 3). It also makes the proof a little cleaner, because every $\frac{1}{5}$-quasirandom group has a unique maximal normal subgroup (Lemma 4.8). We point out that similar methods are useful to study quasirandomnes at other regimes different from the exponential, which could be of interest as well (see also Remark 4.6). The shallow version of Theorem B is the following:
@@
We will actually prove Corollary D as a “baby” Theorem B. The philosophy behind the proofs is that quasirandom groups are pretty rigid, and lend themselves very little to nontrivial extensions, especially the noncentral ones. To make this idea really say something, we use various properties of the finite simple groups and their representations. Particularly for Theorem B, we need a detailed (i.e. quantitative) analysis of quasirandom degrees, modular and projective representations.
The paper is structured as follows. In Section 2 we introduce the (exponential) quasirandom degree $\bm{\varepsilon}(G)$ of a finite group, and give basic results about quasisimple groups, modular and projective representations. In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem Bb. We prove the main theorems in Section 4. Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups, so we collect these data in the Appendix.
@@ 2.1. Minimal degree of a nontrivial representation
We start with the following, which appears as [25, Ex. 1.3.2].
@@ 3. $(iii)$
We would like to mention some connections between $\mathbf{D}_{0}(G)$ and other important parameters associated to a finite group: the number of conjugacy classes $\mathbf{k}(G)$, the representation growth $\mathbf{r}_{n}(G)$, and the minimal index of a proper subgroup, i.e. $\bm{\iota}(G):=\min_{H<G}|G:H|$.
@@
Let $G$ be an $\varepsilon$-quasirandom group. It follows that $\varepsilon<\frac{1}{2}$, and $\mathbf{k}(G)\leq|G|^{1-2\varepsilon}$. Moreover, the number $\mathbf{r}_{n}(G)$ of the inequivalent irreducible representations of degree $n$ is polynomially bounded, namely $\mathbf{r}_{n}(G)\leq n^{\varepsilon^{-1}-2}$ for all $n$ (see [21, Lemma 25]).
@@
for every finite perfect group $G$. A converse inequality, $\mathbf{D}_{0}(G)<\bm{\iota}(G)$, follows from the previous lemmas. Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$). From Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
@@ 2.2. Quasirandom degree
By definition, $\mathbf{D}_{0}(G)=|G|^{\bm{\varepsilon}(G)}$, and $\bm{\varepsilon}(G)$ is the maximum $\varepsilon\in[0,\frac{1}{2})$ such that $G$ is an $\varepsilon$-quasirandom group. Observe that $\bm{\varepsilon}(G)>0$ if and only if $G$ is perfect. We now report the theorem of Nikolov and Pyber mentioned in the Introduction:
@@
Let $N\lhd G$ be a noncentral abelian normal subgroup. Then
@@ 2.3. Quasisimple groups
It is a key property of perfect groups that, if $G$ is perfect, then there is a unique universal perfect central extension of $G$, which we denote by $\widetilde{G}$. This extension is obtained by choosing $N$ isomorphic to the Schur multiplier of $G$, that is, the second homology group $H^{2}(G,\mathbb{Z})$. For more details on this, we refer to [1, Sec. 33].
@@ 2.4. Linear and projective representations
So far, we have only dealt with complex linear representations, but modular and projective representations will be a key tool in the proof of Theorem B. For a prime $r\geq 2$, we use $\mathbf{D}_{r}(G)$ to denote the minimal degree of a nontrivial irreducible modular representation of $G$ over a field of characteristic $r$. We recall that a projective representation of $G$ is a representation of a perfect central extension of $G$. Equivalently, it is a group homomorphism
@@
The degree of the representation is the dimension of $V$. As for linear representations, we denote by $\mathbf{R}_{0}(G)$ the minimal degree of a nontrivial complex projective representation, and by $\mathbf{R}_{r}(G)$ the minimal degree of a nontrivial projective representation over a field of characteristic $r$. We remark that, if $G$ is finite and perfect, and $\widetilde{G}$ is its universal perfect central extension, then
@@
In the remainder of this subsection, we focus on the projective representations of finite simple groups of Lie type. We recall that, if $G$ is a group of Lie type over a field of characteristic $p$, then a modular representation of $G$ over a vector space $V$ is in natural characteristic when the characteristic of $V$ is $p$, while it is in cross-characteristic otherwise.
@@
It follows immediately from the definition. ∎
For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom finite simple groups. The infinite families of $\frac{1}{4}$-quasirandom simple groups are
@@
The relevant informations can be recovered from [15, 24] for classical groups, and from [17, Sec. 4] for exceptional groups (see also the Appendix). We collect these data in Table 1.
@@
Moreover, we point out that the comparison between $\mathbf{R}_{0}(L)$ and $\mathbf{R}_{r}(L)$ ($r\neq p$) can be explicitly found, for all simple groups of Lie type, in [26, Sec. 3].
@@ 3. Quasirandom groups of affine type
In this short section, we provide some highly quasirandom nonquasisimple groups. The key tool is Lemma 3.2 below. Before giving the results, we revise some notions from the representation theory of finite groups. Let $V$ be a finite vector space, and let $V^{*}$ be its dual space. We remark that $V^{*}\cong Hom(V,\mathbb{C}^{*})$. If $\phi:V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}:V^{*}\rightarrow V^{*}$ is defined by
@@
for every $g\in K$. It is well known that the representations $\rho$ and $\rho^{*}$ are not isomorphic in general.
@@
Let $V=(\mathbb{F}_{q})^{n}$ and $g\in K$. The number of $g$-fixed points in $V$ is $q^{\dim(Ker(g-I))}$, while the number of $g$-fixed points in $V^{*}$ is $q^{\dim(Ker((g^{-1})^{T}-I))}$, where $g^{T}$ is the transpose of $g$. Now $\dim(Ker(g-I))$ is the multiplicity of $1$ as an eigenvalue of $g$. We remark that the eigenvalues of $(g^{-1})^{T}$ are the eigenvalues of $g^{-1}$, which are the inverses of the eigenvalues of $g$. In particular, the two quantities above are equal. Since $g$ is arbitrary, the proof follows from Burnside’s lemma, i.e. the number of orbits is equal to the average number of fixed points. ∎
@@ 4.1. Proof of Theorem A
We immediately prove the second part of the theorem. We restate it as follows.
Lemma 4.1.
Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group. Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $|G/Sol(G)|\geq|G|^{\delta}$. If $|G|$ is sufficiently large with respect to $\varepsilon$ and $\delta$, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
Proof.
By [8, Lemma 4.10], we have
$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
[…]
and passing to the logarithms,
The proof follows. ∎
The main idea behind the other direction of Theorem A is the following. Under the hypothesis that $Sol(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$. If this is not the case, then $G/F^{*}(G)$ has a nontrivial simple quotient $L$. We use the quasirandomness of $L$ to obtain an upper bound on $|G|$, while Bender’s theorem and other considerations provide a lower bound on $|F^{*}(G)|$. These inequalities fight each other, giving the desired contradiction. A similar argument will be used in a portion of the proof of Theorem B. Going into details, we start reporting the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
@@
In the following remark, we discuss the asymptotic behaviour of $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$, when $L$ is a finite simple group, and $|L|\rightarrow+\infty$. Of course, we do not care about the sporadic simple groups in this context.
@@
After Lemma 4.1, we have only to prove the first part of the theorem. Fix $\varepsilon>0$. From Corollary 2.3(i), it is clear that $|G/Sol(G)|\geq|G|^{\varepsilon}$. Moreover, $G/Sol(G)$ is again $\varepsilon$-quasirandom. Up to replacing $G$ with $G/Sol(G)$, we can assume that $G$ has a trivial solvable radical. Now it is well known that the generalized Fitting subgroup $F=F^{*}(G)$ coincides with the socle of $G$, and it is the direct product of nonabelian simple groups $T_{i}$’s. Let $F=\prod_{i}(T_{i})^{n_{i}}$, where $n_{i}\geq 1$ for each $i$. Moreover, as $\mathbf{C}_{G}(F)\subseteq F$ [1, (31.13)], we have $\mathbf{C}_{G}(F)=\mathbf{Z}(F)=1$. It follows that $G$ can be embedded as a subgroup of
@@
In particular, all the composition factors of $G/F$ lie as sections of one among $\mathrm{Out}(T_{i})$ or $\mathrm{Sym}(n_{i})$, for some $i$. We claim that $G=F$. Otherwise, it is easy to see that $G/F$ has a simple quotient $L$ which is $\varepsilon$-quasirandom. By the definition of quasirandom degree, we have
(4.1) $|G|\>\leq\>\mathbf{D}_{0}(L)^{1/\varepsilon}.$
So, if $\mathbf{D}_{0}(L)$ is bounded, then $|G|$ is bounded. Let $n=\max_{i}n_{i}$, so that $|F|\geq 60^{n}$. From Remark 4.3 and Theorem 4.2, $L$ has to appear as a section of $\mathrm{Alt}(n)$. From Lemma 4.4, $\mathrm{Sym}(n)$ has a subgroup isomorphic to $L$. From Lemma 2.2 we obtain $\mathbf{D}_{0}(L)\leq n$, and so
@@
When $\mathbf{D}_{0}(L)$ is sufficiently large, this gives a contradiction to (4.1). We have proven that $G/Sol(G)=F^{*}(G/Sol(G))$, so $G/Sol(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/Sol(G)=\prod_{i}(T_{i})^{n_{i}}$. From Remark 4.5, the $T_{i}$’s are all simple groups of Lie type. Let $N=\sum_{i}n_{i}$. From Corollary 2.9, we have $\varepsilon\leq\mathbf{D}_{0}(G/Sol(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$. Moreover, each block $(T_{i})^{n_{i}}$ has to be $\varepsilon$-quasirandom as a quotient of $G$. From Corollary 2.9 and Remark 4.5, we have
@@
Remark 4.6.
An attentive reader will have noticed that a bound of the type $\mathbf{D}_{0}(G)\geq c\log|G|$ is strong enough to give that $G/Sol(G)$ is the direct product of nonabelian finite simple groups.
It is natural to ask whether the condition $|G/Sol(G)|\geq|G|^{\varepsilon}$ is actually superflous in the second part of Theorem A, as it may be implied by the quasirandomness of the sole $G/Sol(G)$. The answer is no, as it can even happen that $|G/Sol(G)|$ is bounded: it is enough to consider the perfect deleted permutation module $(C_{n})^{4}\rtimes\mathrm{Alt}(5)$ when $n$ is large. The following example, where $|G/Sol(G)|$ is unbounded but too small, is due to S. Eberhard.
@@
where $g\in\mathrm{SL}_{2}(p)$, and * is arbitrary. We have
@@
It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix. In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $0.34$-quasirandom for large $p$, but its cardinality is less than $p^{3}$.